OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. m = -4, point (2, -5)

OpenStudy (anonymous):

Use the point-slope formula: -5=(-4)(2)+b -5=-8+b b=-13 y=-4x-13

OpenStudy (anonymous):

Alright, using the formula I used before, y = mx + b, plug in what you know. In this case, using the slope m = -4, we get: y = -4x + b Now, use the point you have (x, y) - > (2, -5), and plug it into the equation. -5 = -8 + b. Solve for b by adding 8 to both sides, and b = 3. Now, go back to the beginning function I wrote y = mx + b. In a linear equation, the slope and y intercept never change. So you write the equation with both those values plugged in. y = -4x + 3. Or, y = 3 - 4x

OpenStudy (anonymous):

Sniping, you have a miscalculation in your third line. It should be b = 3, since -5 + 8 = 3